### v6

@@ 3.2. Extended Transformation
flushleft $\Gamma$Uparrow $\Gamma^{\prime}$
@@
To define repair, the user must provide custom equivalence relations for any type they wish to consider as a setoid. This is done by providing three terms:
@@
fun (x : unit) (H : @eq unit x tt) =>
@eq_ind_r unit tt
(fun x0 : unit => @eq unit x0 tt)
(reflexivity(@eq unit) tt) x H
@@
(reflexivity(eq_unit_two) one), x, H)]]
: forall x : unit_two,
@@
Figure 6. A demonstration of repairing a rewrite in a proof. The left side is over the source type, which is simply the type with one element. We provide a sample Coq tactic script which performs a rewrite, and below it is the term produced. The right side shows the repaired term over the target setoid, which is the type with two elements which are equal under the equivalence relation.
@@ 4. Implementation
We implement this extension to the transformation by extending the Pumpkin Pi Coq plugin (Section 4.1). Consistently with the original Pumpkin Pi, our implementation relies on some user annotations (Section 4.2), and places some other restrictions on the format terms can take (Section 4.3). Our implementation includes custom automation to dispatch proper proofs specific to setoids (Section 4.4).
@@ 4.1. Extending Pumpkin Pi
We extend the Pumpkin Pi Coq plugin directly. Plugins are a method of adding functionality to Coq. They are written in OCaml and can interact directly with the Coq internal codebase. Plugins can directly transform and produce terms; all terms that plugins produce are checked by Coq’s type checker, and so cannot be ill typed.
Pumpkin Pi has various classes of proof repair transformations across type equivalences for which it has specialized automation. We add an additional class, termed setoid lifting, to support our extended transformation. This class mostly reuses the existing transformation, but implements the new rules from Figure 5. Once the transformation rules were determined, the implementation was made straightforward thanks to Pumpkin Pi’s extensibility; in all, our extension adds 1659 lines of code, 510 of which are dedicated to proper proof generation implemented in propergen.ml. The core extensions to the transformation can be found in lift.ml, liftconfig.ml, and liftrules.ml in the supplementary material.
In our extended configuration, the proof that $\equiv_{C}$ is an equivalence relation takes the form of instances of type classes Equivalence $\equiv_{C}$. This makes it possible to use Coq’s setoid automation to implement the oracle that produces proofs of rewrites described in Section 3.2 with no axioms or other extensions to the trusted computing base. Coq has a tactic, called setoid_rewrite, which attempts to perform rewriting by an equivalence relation. However, because equivalence relations are not generally Leibniz, we must prove for each function we define that the function is proper, as defined in Section 2, if we wish to rewrite under applications of that function. In Coq, we prove this by creating an instance of the Proper type class. The setoid_rewrite tactic uses the Proper and Equivalence type class instances to search for proofs of rewrites, and thus we can use it as our oracle.
@@ 4.2. User Annotations
To perform repair, Pumpkin Pi requires that users annotate their proofs explicitly with components of the configuration. These annotations are required to identify parts of the configuration, thereby decoupling the undecidable part of proof repair (configuration inference) from the decidable part (the proof term transformation itself). We inherit this requriement for our extension, and thus both the original configuration and to our extension to equivalences must be annotated. However, we have implemented automation into Pumpkin Pi that generates many of the needed annotations for equivalence relations automatically.
Specifically, if no relation is provided for a given type, our extension defaults to strict equality. Then, we can automatically infer annotations corresponding to applications of equality @eq C, as well as applications of reflexivity @eq_refl C. The same holds for rewrites that fully apply any of Coq’s equality eliminators. However, this annotation inference cannot deal with unapplied instances of @eq and @eq_refl that may later be specialized to C, or equality eliminators that are not fully applied. Such terms are considered improperly annotated.
When an equivalence relation is provided for a type, we provide a custom tactic rewrite_annotate which the user can use in place of rewrite in proofs, which automatically performs annotation while rewriting. This inserts a custom annotation term START_REWRITE before applying the rewrite, which our extension looks for to identify and repair rewrites. This constant takes two non-implicit arguments: the proof of equivalence by which the user is rewriting, and the type they are rewriting.
@@ 4.3. Restrictions
Like the original Pumpkin Pi, we repair proof terms directly, and we do not directly repair terms involving pattern matching and recursion. Pumpkin Pi includes some automation to transform pattern matching and recursion to induction, which is likewise bundled in our extension. We inherit Pumpkin Pi’s proof term decompiler, which makes it possible to get simple tactic proofs from repaired proof terms.
@@
While the original proof repair work had a single depElim term on each side of the configuration, we have multiple eliminators. One eliminates into the sort Type, but is purely nondependent. The other is dependent, but only eliminates into the sort Prop, and requires that the motive of the eliminator be proven to be proper, considering Prop as a setoid with iff. The types of two of these eliminators for one of our case studies are in Figure 7. We do this because Coq’s setoid automation does not work for a dependently typed notion of an equivalence relation. Thus, we cannot use the setoid automation to perform rewrites on applications of functions we define with a dependently typed eliminator. For our Prop-sorted eliminator, this loss means that users cannot automatically perform rewrites on the proofs of propositions. In Coq, Prop is frequently treated as effectively proof irrelevant, so this loss is more acceptable.
@@
In addition, our use of Coq’s setoid automation is facilitated through the use of Coq’s rewrite tactics. However, these tactics do not allow a rewrite which does not change the goal type, and so neither does our automation for repairing setoid rewrites. Furthermore, the setoid rewrite automation does not allow specifying a motive P. To perform setoid rewrites with a motive, a user trying to prove P x -> P y can perform a substitution P[z/y], where z is free in P, and then define Q := fun z => P[z/y]. Then, for another fresh variable w, the user can perform a setoid_rewrite to prove H : $\forall$ (w : B), Q w x -> Q w y, and recover the desired rewrite proof as H y. Our extension uses this methodology when repairing rewrites with a motive to setoids.
4.4. Automating Proper Proofs
@@
To perform rewrites on a term using Coq’s setoid automation, it is necessary to prove that the functions in that term are proper. Thus, when repairing terms, it is potentially necessary to have proper proofs for every previously repaired function. We implement automation that helps prove many functions to be proper automatically. Our automation must fail in some cases, though, since proving a general function is proper is undecidable. To prove this, for any proposition P : Prop, define f : bool -> Prop, f b = if b then True else P. Generating a proof that f is proper, using bool as a setoid relating true and false, and Prop as a setoid with the relation iff, is equivalent to proving iff P True, which is undecidable.
@@
Presently, our automation constructs proper proofs automatically in two practical cases. First, when f is a composition of proper functions, we introduce hypotheses stating that all arguments are equivalent, and then rewrite by these hypotheses. This is the approach taken by Coq’s
@@
solve_proper tactic, though we modify it slightly. Coq’s solve_proper fails if any of its inputs do not appear in the body of the function because the rewrite for that argument will fail. To avoid this, we use the try tactical when rewriting. We also try both Coq’s rewrite and setoid_rewrite tactics, while solve_proper only runs setoid_rewrite. While setoid_rewrite supports rewriting under binders in some instances, rewrite succeeds in some instances where setoid_rewrite fails.
@@
This can be done in general for inductive types. Thus, our automation checks if f is an eliminator at the top level, and if so generates and tries to prove such a proper goal. If it succeeds, the automation attempts to show that the inductive cases (in the above example, ps) are proper using the rewriting strategy described above. In our case studies, this is necessary for automatically solving some of the generated proper goals.
@@
Figure 8. The types of our integer representations. We provide an instance of Equivalence eq_GZ for the case study in the supplementary material.
@@ 5. Case Studies
We use our extended version of Pumpkin Pi to automatically repair proofs on three case studies that use quotient type equivalences. First, we conduct repair between two representations of the integers (Section 5.1). Second, we study two common implementations of the queue data structure and how we can repair from one to the other (Section 5.2). Third, we repair between dense and sparse representations of polynomials with natural number coefficients (Section 5.3). In each of these case studies, the type to which we repair functions and theorems has a structure which enables efficient implementations of key operations. Our extension makes it possible to take advantage of this structure while using the repaired theorems, something impossible with prior versions of Pumpkin Pi. All of the case study examples can be found in more detail in the supplementary material.
@@ 5.1. Adding, Fast and Slow
Our first case study is a usage of quotients that is foundational in mathematics. We consider a change in the type representing integers from the inductive type found in many standard libraries to the standard quotient based representation. We repair addition and proofs about addition from one representation to the other. Finally, we recover the repaired proofs for a more efficient version of addition over the repaired type. Our Coq implementation of this case study can be found in grothendieck_int_equivalence_repair_tool.v.
@@
Our first representation, Z, is based on the default implementation of the integers in Cubical Agda: two copies of $\mathbb{N}$ glued back-to-back. We will repair functions and proofs about Z to use a representation that may be more familiar to set theorists: viewing the integers as elements of $\mathbb{N}\times\mathbb{N}/\sim$, where $(x_{1},x_{2})\sim(y_{1},y_{2})\iff x_{1}+y_{2}=x_{2}+y_{1}$. We call the resulting type GZ [^3] In reference to Grothendieck, as this is the Grothendieck group of the natural numbers, and the equivalence relation eq_GZ. The definitions are in Figure 8. Z and GZ are setoid equivalent using the definition from Section 2, by mapping pos n to (n, 0) and negsuc n to (0, S n).
We next decompose our isomorphism into a repair configuration consisting of dependent constructors, dependent eliminators, and $\iota$-reduction rules for both types. The configuration differs from those found in the original Pumpkin Pi examples in that there are two eliminators (Figure 7 from Section 4.3): depRec for eliminating into nondependent types, and depElimProp for eliminating into dependent types that reside in Prop. Furthermore, depElimProp on GZ has an extra proof obligation: the motive P : GZ -> Prop must be a proper function, where the sort Prop is viewed as the setoid (Prop, iff). While in theory each of these eliminators need their own set of $\iota$-reduction theorems, we provide them solely for depRec, since needing $\iota$ for depElimProp is not common and does not show up in our case study. The full configuration can be found in the supplementary material.
@@
Our extension of Pumpkin Pi also automatically generates proofs that the repaired functions are proper. First, the user must prove that depRec is proper. Our extension uses this to generate proper proofs for functions applying depRec. For example, it generates the proof that addition is proper:
@@
We automatically repair the proof add0LZ, which shows that 00 is a left identity for addition. Figure 10 shows the old and new theorem types. Note that, in the repaired theorem type, equality has been automatically replaced with an equivalence relation on the type, reflecting that the repaired theorem is about setoid equality instead of eq. In addition, we repair a proof that 00 is a right identity for addition. The original and repaired proofs can be found in the supplementary material.
For now, there is a bit of extra work related to proper proof generation when our proofs apply depElimPropGZ (as add0LZ does). In particular, while our automation generates proper proofs for some of the motives passed to depElimPropGZ, there is not yet a way to automatically supply those proofs to Pumpkin Pi so that it uses them when repairing applications of depElimPropZ. For now, we define a constant corresponding to the motive, which we then separately repair:
@@
appliedDepElimPropZ before reconfiguring. Future versions can avoid the need for this workaround by automatically supplying the necessary proper proofs to depElimProp terms.
@@
Figure 12. One list queues and two list queues. We provide an instance of Equivalence eq_queue in the supplementary material.
@@ 5.2. Variations on a Theme of Queues
Next, we repair functions and proofs across a change in implementation of a queue data structure. This is motivated by an example from Angiuli et al. (2021), which showed that quotient types can be used to adjust certain relations more general than equivalences into equivalences for use with transport in Cubical Agda. That class of changes was cited in the Pumpkin Pi paper as an example that could not be expressed naturally in Coq with the original framework. With our extensions to Pumpkin Pi, we can express this using setoids. Our Coq implementation of this case study can be found in two_list_queue_equivalence_repair_tool.v.
@@
Each two list queue (l1, l2) corresponds to the queue l1 ++ (rev l2), but multiple two list queues correspond to a single one list queue. Thus, we use the equivalence relation (l1, l2) $\sim$ (l3, l4) $\iff$ l1 ++ (r%ev l2) = l3 ++ (rev l4) and consider TLQ as a setoid. These two types are setoid equivalent along the expected correspondence, which lets us define the repair configuration. See Figure 12 for the types, and the supplementary material for the repair configuration.
@@
In the previous case study, our automation succeeded in generating every function’s proper proof. This time, however, while the proofs that enqueueTLQ and dequeueTLQ are proper are generated automatically, the proper proofs for dequeueHelpTLQ and returnOrEnqTLQ failed to generate and needed to be supplied manually. Also, we need to define multiple equivalence relations, since the return type of dequeueTLQ is option (TLQ * A) and our automation does not yet automatically lift the equivalence over TLQ to types including TLQ. The user must provide these equivalences to Pumpkin Pi.
@@
Figure 16. Definitions and equivalence relations for CLPoly and CEPPoly. removeLeadingZeros l removes leading zeros from l, while coeff p exp is the nth degree coefficient of p. We provide instances of Equivalence eq_CLPoly and Equivalence eq_CEPPoly in the supplementary material.

### v7

@@ 3.2. Extended Transformation
flushleft $\Gamma$ $\Uparrow$ $\Gamma^{\prime}$
@@
To define an instance of a repair transformation using setoids, the user must provide custom equivalence relations for any type they wish to consider as a setoid. This is done by providing three terms:
@@
fun (x : unit) (H : @eq unit x $t$) =>
@eq_ind_r unit $t$
(fun x0 : unit => @eq unit x0 $t$)
(reflexivity(@eq unit) $t$) x H
@@
(reflexivity(eq_unit_two) one), x, H)
]] : forall x : unit_two,
@@
Figure 6. A demonstration of repairing a rewrite in a proof. The left side is over the source type, which is simply the type with one element. We provide a sample Rocq tactic script which performs a rewrite, and below it is the term produced. The right side shows the repaired term over the target setoid, which is the type with two elements which are equal under the equivalence relation.
@@ 4. Implementation
We implement this extension to the transformation by extending the Pumpkin Pi Rocq plugin (Section 4.1). Consistently with the original Pumpkin Pi, our implementation relies on some user annotations (Section 4.2), and places some other restrictions on the format terms can take (Section 4.3). Our implementation includes custom automation to dispatch properness proofs specific to setoids (Section 4.4).
@@ 4.1. Extending Pumpkin Pi
We extend the Pumpkin Pi Rocq plugin directly. Plugins are a method of adding functionality to Rocq. They are written in OCaml and can interact directly with the Rocq internal codebase. Plugins can directly transform and produce terms; all terms that plugins produce are checked by Rocq’s type checker, and so cannot be ill typed.
Pumpkin Pi has various classes of proof repair transformations across type equivalences for which it has specialized automation. We add an additional class, termed setoid lifting, to support our extended transformation. This class mostly reuses the existing transformation, but implements the new rules from Figure 5. Once the transformation rules were determined, the implementation was made straightforward thanks to Pumpkin Pi’s extensibility; in all, our extension adds 1659 lines of code, 510 of which are dedicated to properness proof generation implemented in propergen.ml. The core extensions to the transformation can be found in lift.ml,   4 liftconfig.ml,   5 and liftrules.ml   6 in the artifact.
In our extended configuration, the proof that $\equiv_{C}$ is an equivalence relation takes the form of instances of type classes Equivalence $\equiv_{C}$. This makes it possible to use Rocq’s setoid automation to implement the oracle that produces proofs of rewrites described in Section 3.2 with no axioms or other extensions to the trusted computing base. Rocq has a tactic, called setoid_rewrite, which attempts to perform rewriting by an equivalence relation. However, because rewriting arbitrary terms between equivalent elements is not possible, we must prove for each function we define that the function is proper, as defined in Section 2, if we wish to rewrite under applications of that function. In Rocq, we prove this by creating an instance of the Proper type class. The setoid_rewrite tactic uses the Proper and Equivalence type class instances to search for proofs of rewrites, and thus we can use it as our oracle.
@@ 4.2. User Annotations
To perform repair, Pumpkin Pi requires that users annotate their proofs explicitly with components of the configuration. These annotations are required to identify parts of the configuration, thereby decoupling the undecidable part of proof repair (configuration inference) from the decidable part (the proof term transformation itself). We inherit this requirement for our extension, and thus both the original configuration and to our extension to equivalences must be annotated. However, we have implemented automation into Pumpkin Pi that generates many of the needed annotations for equivalence relations automatically.
Specifically, if no relation is provided for a given type, our extension defaults to strict equality. Then, we can automatically infer annotations corresponding to applications of equality @eq C, as well as applications of reflexivity @eq_refl C. The same holds for rewrites that fully apply any of Rocq’s equality eliminators   5. However, this annotation inference cannot deal with unapplied instances of @eq and @eq_refl that may later be specialized to C, or equality eliminators that are not fully applied. Such terms are considered improperly annotated.
When an equivalence relation is provided for a type, we provide a custom tactic rewrite_annotate which the user can use in place of rewrite in proofs, which automatically performs annotation while rewriting. This inserts a custom annotation term START_REWRITE before applying the rewrite, which our extension looks for to identify and repair rewrites. This constant takes two non-implicit arguments: the proof of equivalence by which the user is rewriting, and the type they are rewriting   7.
@@ 4.3. Restrictions
Definition respectful_hetero
(A B : Type)
(C : A -> Type) (D : B -> Type)
(R : A -> B -> Prop)
(R : forall (x : A) (y : B), C x -> D y -> Prop) :
(forall x : A, C x) -> (forall x : B, D x) -> Prop :=
fun f g => forall x y, R x y -> R x y (f x) (g y).
Figure 8. An equivalence relation from the Rocq standard library on dependently typed functions stating that the functions are respectful. R relates elements of their domains, and R relates elements of their codomains. Notice that because the functions are dependently typed, R is a family of equivalence relations, one for each pair of elements x : A, y : B. Specializing respectful_hetero to non-dependently typed functions is used to define the notion of a proper function we use in this work.
@@
Pumpkin Pi does not directly repair terms involving pattern matching and recursion, instead requiring that proof terms it repairs are written using induction principles. As a result, our extension to Pumpkin Pi also has this limitation. However, Pumpkin Pi includes some automation to transform pattern matching and recursion in proof terms into induction, which is likewise bundled in our extension. We inherit Pumpkin Pi’s proof term decompiler, which makes it possible to get simple tactic proofs from repaired proof terms.
@@
While the original proof repair work had a single depElim term on each side of the configuration, we have multiple eliminators. One eliminates into the sort Type, but is purely nondependent. The other is dependent, but only eliminates into the sort Prop, and requires that the motive of the eliminator be proven to be proper, considering Prop as a setoid with iff. The types of two of these eliminators for one of our case studies are in Figure 7. We do this because Rocq’s setoid automation does not work for a dependently typed notion of a proper function. Specifically, the term respectful_hetero, given in Figure 8, exists in the standard library (Team, 2019) and could form the basis for such automation, but no such automation has been implemented in the standard library. Thus, we cannot use the setoid automation to perform rewrites on applications of functions we define with a dependently typed eliminator. For our Prop-sorted eliminator, this loss means that users cannot automatically perform rewrites on the proofs of propositions. In Rocq, Prop is frequently treated as effectively proof irrelevant, so this loss is more acceptable.
@@
In addition, our use of Rocq’s setoid automation is facilitated through the use of Rocq’s rewrite tactics. However, these tactics do not allow a rewrite which does not change the goal type, and so neither does our automation for repairing setoid rewrites. Furthermore, the setoid rewrite automation does not allow specifying a motive P. To perform setoid rewrites with a motive, a user trying to prove P x -> P y can perform a substitution P[z/y], where z is free in P, and then define Q := fun z => P[z/y]. Then, for another fresh variable w, the user can perform a setoid_rewrite to prove H : $\forall$ (w : B), Q w x -> Q w y, and recover the desired rewrite proof as H y. Our extension uses this methodology when repairing rewrites with a motive to setoids   4.
4.4. Automating Properness Proofs
@@
To perform rewrites on a term using Rocq’s setoid automation, it is necessary to prove that the functions in that term are proper. Thus, when repairing terms, it is potentially necessary to have properness proofs for every previously repaired function. We implement automation that helps prove many functions to be proper automatically. Our automation must fail in some cases, though, since proving a general function is proper is undecidable. To prove this, for any proposition P : Prop, define f : bool -> Prop, f b = if b then True else P. Generating a proof that f is proper, using bool as a setoid relating true and false, and Prop as a setoid with the relation iff, is equivalent to proving iff P True, which is undecidable.
@@
Presently, our automation constructs properness proofs automatically in two practical cases. First, when f is a composition of proper functions, we introduce hypotheses stating that all arguments are equivalent, and then rewrite by these hypotheses. This is the approach taken by Rocq’s
@@
solve_proper tactic, though we modify it slightly. Rocq’s solve_proper fails if any of its inputs do not appear in the body of the function because the rewrite for that argument will fail. To avoid this, we use the try tactical when rewriting. We also try both Rocq’s rewrite and setoid_rewrite tactics, while solve_proper only runs setoid_rewrite. While setoid_rewrite supports rewriting under binders in some instances, rewrite succeeds in some instances where setoid_rewrite fails.
@@
This can be done in general for inductive types. Thus, our automation checks if f is an eliminator at the top level, and if so generates and tries to prove such a proper goal. If it succeeds, the automation attempts to show that the inductive cases (in the above example, ps) are proper using the rewriting strategy described above   8. In our case studies, this is necessary for automatically solving some of the generated proper goals   9.
@@
Figure 9. The types of our integer representations. We provide an instance of Equivalence eq_GZ for the case study in the artifact.
@@ 5. Case Studies
We use our extended version of Pumpkin Pi to automatically repair proofs on three case studies that use quotient type equivalences. First, we conduct repair between two representations of the integers (Section 5.1). Second, we study two common implementations of the queue data structure and how we can repair from one to the other (Section 5.2). Third, we repair between dense and sparse representations of polynomials with natural number coefficients (Section 5.3). In each of these case studies, the type to which we repair functions and theorems has a structure which enables efficient implementations of key operations. Our extension makes it possible to take advantage of this structure while using the repaired theorems, something impossible with prior versions of Pumpkin Pi. All of the case study examples can be found in more detail in the artifact.
@@ 5.1. Adding, Fast and Slow
Our first case study is a usage of quotients that is foundational in mathematics. We consider a change in the type representing integers from the inductive type found in many standard libraries to the standard quotient based representation. We repair addition and proofs about addition from one representation to the other. Finally, we recover the repaired proofs for a more efficient version of addition over the repaired type. Our Rocq implementation of this case study can be found in the artifact   9.
@@
Our first representation, Z, is based on the default implementation of the integers in Cubical Agda: two copies of $\mathbb{N}$ glued back-to-back. We will repair functions and proofs about Z to use a representation frequently used as a definition in mathematics: viewing the integers as elements of $\mathbb{N}\times\mathbb{N}/\sim$, where $(x_{1},x_{2})\sim(y_{1},y_{2})\iff x_{1}+y_{2}=x_{2}+y_{1}$. We call the resulting type GZ [^4] In reference to Grothendieck, as this is the Grothendieck group of the natural numbers, and the equivalence relation eq_GZ. The definitions are in Figure 9. Z and GZ are setoid equivalent using the definition from Section 2, by mapping pos n to (n, 0) and negsuc n to (0, S n).
We next decompose our isomorphism into a repair configuration consisting of dependent constructors, dependent eliminators, and $\iota$-reduction rules for both types. The configuration differs from those found in the original Pumpkin Pi examples in that there are two eliminators (Figure 7 from Section 4.3): depRec for eliminating into nondependent types, and depElimProp for eliminating into dependent types that reside in Prop. Furthermore, depElimProp on GZ has an extra proof obligation: the motive P : GZ -> Prop must be a proper function, where the sort Prop is viewed as the setoid (Prop, iff). While in theory each of these eliminators need their own set of $\iota$-reduction theorems, we provide them solely for depRec, since needing $\iota$ for depElimProp is not common and does not show up in our case study. The full configuration can be found in the artifact.
@@
Our extension of Pumpkin Pi also automatically generates proofs that the repaired functions are proper. First, the user must prove that depRec is proper. Our extension uses this to generate properness proofs for functions applying depRec. For example, it generates the proof that addition is proper:
@@
We automatically repair the proof add0LZ, which shows that 0 is a left identity for addition. Figure 11 shows the old and new theorem types. Note that, in the repaired theorem type, equality has been automatically replaced with an equivalence relation on the type, reflecting that the repaired theorem is about setoid equality instead of eq. In addition, we repair a proof that 0 is a right identity for addition. The original and repaired proofs can be found in the artifact.
For now, there is a bit of extra work related to proper proof generation when our proofs apply depElimPropGZ (as add0LZ does). In particular, while our automation generates properness proofs for some of the motives passed to depElimPropGZ, there is not yet a way to automatically supply those proofs to Pumpkin Pi so that it uses them when repairing applications of depElimPropZ. For now, we define a constant corresponding to the motive, which we then separately repair:
@@
appliedDepElimPropZ before reconfiguring. Future versions can avoid the need for this workaround by automatically supplying the necessary properness proofs to depElimProp terms.
@@
Figure 13. One list queues and two list queues. We provide an instance of Equivalence eq_queue in the artifact.
@@ 5.2. Variations on a Theme of Queues
Next, we repair functions and proofs across a change in implementation of a queue data structure. This is motivated by an example from Angiuli et al. (2021), which showed that quotient types can be used to adjust certain relations more general than equivalences into equivalences for use with transport in Cubical Agda. That class of changes was cited in the Pumpkin Pi paper as an example that could not be expressed naturally in Rocq with the original framework. With our extensions to Pumpkin Pi, we can express this using setoids. Our Rocq implementation of this case study can be found in the artifact   10.
@@
Each two list queue (l1, l2) corresponds to the queue l1 ++ (rev l2), but multiple two list queues correspond to a single one list queue. Thus, we use the equivalence relation (l1, l2) $\sim$ (l3, l4) $\iff$ l1 ++ (rev l2) = l3 ++ (rev l4) and consider TLQ as a setoid. These two types are setoid equivalent along the expected correspondence, which lets us define the repair configuration. See Figure 13 for the types, and the artifact for the repair configuration.
@@
In the previous case study, our automation succeeded in generating every function’s properness proof. This time, however, while the proofs that enqueueTLQ and dequeueTLQ are proper are generated automatically, the properness proofs for dequeueHelpTLQ and returnOrEnqTLQ failed to generate and needed to be supplied manually. Also, we need to define multiple equivalence relations, since the return type of dequeueTLQ is option (TLQ * A) and our automation does not yet automatically lift the equivalence over TLQ to types including TLQ. The user must provide these equivalences to Pumpkin Pi.
@@
Figure 17. Definitions and equivalence relations for CLPoly and CEPPoly. removeLeadingZeros l removes leading zeros from l, while coeff p exp is the nth degree coefficient of p. We provide instances of Equivalence eq_CLPoly and Equivalence eq_CEPPoly in the artifact.